OpenStudy (anonymous):

find the derivative of the function. y=(x^2-3x+2)/(x^7-2)

OpenStudy (anonymous):

I did that, but my answer is not like the answer choices.

OpenStudy (anonymous):

sorry my answer was slightly wrong there

OpenStudy (anonymous):

\[\frac{d}{dx}\frac{f(x)}{g(x)} = \frac{f'(x)g(x) - g'(x)f(x)}{(g(x))^2}\]

OpenStudy (anonymous):

\[f(x) = x^2 - 3x + 2\] so\[f'(x) = 2x - 3\] and \[g(x) = x^7 - 2\] so\[g'(x) = 7x^6\]

OpenStudy (anonymous):

ok I did that

OpenStudy (anonymous):

\[\frac{d}{dx} \frac{x^2 - 3x + 2}{x^7 - 2} = \frac{(2x - 3)(x^7 - 2) - 7x^6(x^2 - 3x + 2)}{(x^7 - 2)^2}\]

OpenStudy (anonymous):

now we probably have to see if it can be simplified

OpenStudy (anonymous):

on my answer choices i have to end up with a x^8 i dont see how i can get that

OpenStudy (anonymous):

on the top?

OpenStudy (anonymous):

is it a -5x^8

OpenStudy (anonymous):

yes how did you get that?

OpenStudy (anonymous):

did you get my earlier answer? just multiply out the top and collect like terms

OpenStudy (anonymous):

i did and i still end up with just x^7

OpenStudy (anonymous):

double check your multiplication to make sure you haven't made a silly mistake :)

OpenStudy (anonymous):

ok thank you